# Multi-target quantum compilation algorithm

### v1

@@ Abstract
In quantum computing, quantum compilation involves transforming information from a target unitary into a trainable unitary represented by a quantum circuit. Traditional quantum compilation optimizes circuits for a single target. However, many quantum systems require simultaneous optimization of multiple targets, such as simulating systems with varying parameters and preparing multi-component quantum states. To address this, we develop a multi-target quantum compilation algorithm to enhance the performance and flexibility of simulating multiple quantum systems. Through our benchmarks and case studies, we demonstrate the algorithm’s effectiveness, highlighting the significance of multi-target optimization in the advancement of quantum computing. This work establishes the groundwork for further development, implementation, and evaluation of multi-target quantum compilation algorithms.
keywords:
genetic algorithm, quantum compilation, quantum architecture search, quantum optimization
1 Introduction
Variational quantum algorithms (VQAs) like quantum approximate optimization algorithm (QAOA), variational quantum eigensolver (VQE), quantum neural networks (QNN), and quantum compilation (QC) are promising for solving practical tasks on noisy intermediate scale quantum (NISQ) devices beyond classical computers [1]. Recent achievements demonstrate effectiveness in quantum state preparation [2, 3, 4, 5, 6], quantum dynamic simulation [2, 7, 8, 9], and quantum metrology [10, 11, 12, 13, 14]. QC, in particular, has gained significant interest. It uses a training process to transform information from an unknown target unitary into a trainable known unitary [15, 16]. This method has various applications, including gate optimization [15], quantum-assisted compiling [16], continuous-variable quantum learning [17], quantum state tomography [18], and quantum objects simulation [2]. For instance, a quantum object such as a quantum state can be prepared and its evolution can be simulated in a quantum circuit by using QC [2].
@@ 1 Introduction
The performance of QCs relies on the number of qubits and the circuit depth. The choice of ansatzes (trainable quantum circuits) is also crucial and must be carefully selected. Some entangled topologies have shown promise in solving quantum state preparation problems but require many resources due to their numerous layers and parameters [6]. Traditional QCs focus on compiling a single target, like the topography of a single quantum state [18] or the preparation of a single quantum object (e.g., a quantum state, Hamiltonian) [2]. However, many scenarios require multiple targets, such as simulating systems with varying parameters, preparing multi-component quantum states, tomography for multiple states, and optimizing algorithms for different instances. Therefore, it is crucial to develop an effective automatic approach for these tasks.
@@
This work presents a multi-target quantum compilation algorithm to address the need to handle multiple targets. We compile multiple unitaries into a single trainable unitary through an optimization process, improving QC performance and applicability in practical scenarios. We explain the theoretical foundations of our algorithm, its structure, and using techniques like the genetic algorithm (GA) to enhance the multi-target QC process. We show the effectiveness of our algorithm through benchmarking and case studies, including the preparation of thermal states and the simulation of time-dependent Hamiltonians. The results highlight the importance of multi-target compilation in advancing quantum computing technologies and the potential for further innovation in this field.
2 Results
2.1 Multi-target quantum compilation
@@
Quantum compilation (QC) [16, 6] is a training process that transforms information from a target unitary $U$ to a trainable unitary $V(\bm{\theta})$, where $\bm{\theta}$ are trainable parameters. Hereafter, we introduce a compilation process to transform a set of $n$ targets into one trainable unitary, therefore, it is called multi-target quantum compilation.
Definition
@@
where $\bm{\theta}_{j}\in\big{\{}\theta_{j}^{(1)},\theta_{j}^{(2)},...,\theta_{j}^{(m)}\big{\}}$ is a set of $m$ trainable parameters and $\phi$ is a globe phase. For each $U_{j}$, we need to maximize the kernel $\mathcal{K}_{U_{j},V}(\bm{\theta}_{j})=\left|\left\langle\psi\left|U_{j}V^{\dagger}(\bm{\theta}_{j})\right|\psi\right\rangle\right|^{2}$ where $|\psi\rangle$ is a reference state. For maximizing all $\mathcal{K}_{U_{j},V}(\bm{\theta}_{j}),\forall U_{j}\in\mathbf{\mathcal{U}}$, we define a cost function $\mathcal{L}$, which is the average infidelity
$\mathcal{L}(\bm{\Theta})=1-\frac{1}{n}\sum_{j=1}^{n}\mathcal{K}_{U_{j},V}(\bm{\theta}_{j}),$ (2)
where $\bm{\Theta}=(\bm{\theta}_{1},\cdots,\bm{\theta}_{n})$ are trainable parameters, and $0\leq\mathcal{L}(\bm{\Theta})\leq 1$. The optimization process is to find $\bm{\Theta}^{*}=\operatorname*{argmin}_{\{\bm{\Theta}\}}\mathcal{L}(\bm{\Theta}).$
Quantum compilation algorithm
@@
where $G_{k}(\bm{\theta}_{k})$ is a sequence of parameterized single- and multi-qubit quantum gates, $L$ is the number of layers, $N$ the number of qubits. The design of $V(\bm{\theta})$ is not unique, and recent efforts to optimize it, including architecture search [19, 20, 21, 22], adaptive variational quantum algorithm [23, 7], and genetic-based approach (GA) [24, 25, 26, 27].
In this work, we combine GA with VQA to optimize multi-target quantum compilation. This approach optimizes both the parameters and the structure of quantum circuits, ensuring efficient compilation, practical computational capability, and low-depth circuits. The scheme is shown in Fig.1a, and the flow chart is in Fig.1b. The genetic search identifies the best ansatz $V^{*}$ that meets all inputs at a given threshold. The best ansatz is then used in VQA to find the optimal parameters $\bm{\Theta}^{*}$ using gradient descent.
Denote the set of generated quantum circuits as $\mathbf{\mathcal{V}}=\big{\{}V_{l};1\leq l\leq n_{\mathbf{\mathcal{V}}}\big{\}}$, where $n_{\mathbf{\mathcal{V}}}$ is the number of circuits. Each circuit $V_{l}\in\mathbf{\mathcal{V}}$ is created from a pool of quantum gates. For each circuit $V_{l}$, we evaluate the average infidelity using
$\mathcal{L}(\bm{\Theta},V_{l})=1-\frac{1}{n}\sum_{j=1}^{n}\mathcal{K}_{U_{j},V_{l}}(\bm{\theta}_{j})\;,\forall 1\leq l\leq n_{\mathbf{\mathcal{V}}}.$ (4)
Following the flowchart, if $\mathcal{L}(\bm{\Theta},V_{l})\leq\rm threshold$, we designate $V_{l}$ to the best circuit, $V^{*}\leftarrow V_{l}$, and $\bm{\Theta}$ will be the optimal parameters, $\bm{\Theta}^{*}\leftarrow\bm{\Theta}$. If not, we run the VQA scheme with gradient descent to update $\bm{\Theta}$ for a certain number of iterations $n_{\rm iter}$. Then, we use GA with selection, crossover, and mutation to create the next generation of $\mathcal{V}$. See the Method for a detailed evolution of the GA. The GA scheme repeats until the threshold is met or the number of generations $n_{\rm gene}$ is reached. Finally, if $n_{\rm gene}$ is reached without meeting the threshold, we pass the best circuit $V^{*}$ to the VQA scheme again, and optimize $\bm{\Theta}$ until the threshold is met.
@@
Our scheme differs from the one-stage [19] and two-stage architecture search [21, 28]. In those methods, all circuits are generated, their cost functions are evaluated, and they are sorted from best to worst to select the top one. In our approach, each circuit $V_{l}\in\mathcal{V}$ is evaluated and compared with the best circuit (initially set as $V_{1}$). We thus do not require space to store all circuits. Although genetic algorithms have been suggested for quantum state tomography [29], we believe they are also promising for different types of VQAs, including our proposed method for quantum compilation.
Figure 1: Integrated Genetic Approach with Variational Quantum Algorithm (GA-VQA) for multi-target quantum compilation optimization. (a) The GA-VQA scheme involves a parameterized quantum circuit and a classical computer to evaluate and update the circuit and its parameters. (b) The flowchart: (i) Generate a set of circuits $\mathcal{V}$ and parameters $\bm{\Theta}$, (ii) Evaluate the cost function $\mathcal{L}(\bm{\Theta},V_{l})$ for all $l$. (iii) If it fails to pass, run VQA followed by GA to find the best circuit, checking the threshold each time. (iv) If the threshold is not met, pass the best circuit $V^{*}$ to VQA and repeat until the threshold is met.
Numerical illustration
@@
For numerical evaluation, we follow Ref. [30] and divide a random set $\mathcal{U}$ into a training set $\mathcal{U}_{\text{train}}=\{U_{1},U_{2},...,U_{\rm train}\}$ and a testing set $\mathcal{U}_{\text{test}}=\{U_{{\rm train+1}},U_{\rm train+2},...,U_{n}\}$. We then train the model on $\mathcal{U}_{\text{train}}$ using the cost function $\mathcal{L}$, where
$\mathcal{L}(\bm{\Theta}_{\rm train},V_{l})=1-\frac{1}{n_{\rm train}}\sum_{j=1}^{n_{\rm train}}\mathcal{K}_{U_{j},V_{l}}(\bm{\theta}_{j}),$ (5)
After training, we get the best circuit $V^{*}$ and evaluate the expected risk $\mathcal{R}$
@@
$\mathcal{R}(\bm{\Theta}_{\rm test})=\frac{1}{4}\mathbb{E}_{U_{j}\;\in\;\mathcal{U}_{\text{test}}}\left\|U_{j}|\bm{0}\rangle\langle\bm{0}|U_{j}^{\dagger}-[V^{*}]^{\dagger}(\bm{\theta}_{j})|\bm{0}\rangle\langle\bm{0}|V^{*}(\bm{\theta}_{j})\right\|_{1}^{2}.$ (6)
Note that we fixed $|\psi\rangle=|\bm{0}\rangle\equiv|0\rangle^{\otimes N}$.
@@
Figure 2: Numerical benchmarking for a set of Haar random $\mathcal{U}$. (a-d) Correlation between risk and fidelity for different number of generations $n_{\rm gene}$ and circuits per generation $n_{\mathcal{V}}$. (e) Plot of fidelity versus depth $d$ for different number of qubits $N$ ranging from 2 to 5. (f) The best quantum circuit generated by GA-VQA. (g) The quantum circuit generated by Qiskit. (h) $|g_{2}g_{N}\rangle$ (2 layers) ansatz before transforming into single and two-qubit gates.
@@
Figure 2(a-d) shows the correlation between risk $\mathcal{R}$ and fidelity $\mathcal{F}=1-\mathcal{L}$ for several $n_{\rm gene}$ and $n_{\mathcal{V}}$ with $N=3$. The GA-VQA method identifies a low-risk, high-fidelity ansatz for multi-target Haar random unitaries. Additionally, the factors $n_{\rm gene}$ and $n_{\mathcal{V}}$ have minimal impact on the results, allowing GA-VQA to run efficiently with minimal values, thus saving computational resources. Figure 2e shows fidelity versus depth $d$ for 2-5 qubits. As expected, higher $N$ requires larger $d$ to achieve high fidelity. In the figure, circles mark the depth where $\mathcal{F}\geq 0.99$, with upper bounds generated by the standard Qiskit method. Figure 2f shows the best circuit generated by our GA-VQA compared to the Qiskit-generated circuit in Fig. 2g and the $|g_{2}g_{N}\rangle$ ansatz[6] in Fig. 2h for $N=3$. The GA-VQA circuit has the smallest depth. Table 1 shows detailed comparisons for 2 to 5 qubits, demonstrating that the GA-VQA circuit provides the highest fidelity with the smallest depth.
@@
2.2 Applications
Thermal state preparation (TSP)

### v2

@@ Abstract
Quantum compilation is the process of converting a target unitary operation into a trainable unitary represented by a quantum circuit. It has a wide range of applications, including gate optimization, quantum-assisted compiling, quantum state preparation, and quantum dynamic simulation. Traditional quantum compilation usually optimizes circuits for a single target. However, many quantum systems require simultaneous optimization of multiple targets, such as thermal state preparation, time-dependent dynamic simulation, and others. To address this, we develop a multi-target quantum compilation algorithm to improve the performance and flexibility of simulating multiple quantum systems. Our benchmarks and case studies demonstrate the effectiveness of the algorithm, highlighting the importance of multi-target optimization in advancing quantum computing. This work lays the groundwork for further development and evaluation of multi-target quantum compilation algorithms.
I Introduction
Variational quantum algorithms (VQAs) like quantum approximate optimization algorithm (QAOA), variational quantum eigensolver (VQE), quantum neural networks (QNN), and quantum compilation (QC) are promising for solving practical tasks on noisy intermediate scale quantum (NISQ) devices beyond classical computers Cerezo2021 . VQAs have proven to be versatile in various applications, including quantum state preparation Kuzmin2020variationalquantum ; Sagastizabal2021 ; PhysRevLett.129.230504 ; castro2024variational ; hai2023variational ; HAI2024101726 , quantum dynamic simulation HAI2024101726 ; PRXQuantum.2.030307 ; Luo2024 ; PhysRevResearch.6.023130 , and quantum metrology Koczor_2020 ; 9605341 ; Meyer2021 ; Le2023 ; Cimini2024 . Recent developments underscore their ability to tackle complex quantum systems effectively.
In quantum state preparation, significant progress has led to various effective methods for generating target states with high fidelity Kuzmin2020variationalquantum ; Sagastizabal2021 ; PhysRevLett.129.230504 ; castro2024variational ; hai2023variational ; HAI2024101726 . For instance, Zhang et al. demonstrated that any quantum state can be prepared with a linear-depth circuit using many ancillary qubits, leading to exponential speedups for tasks like Hamiltonian simulation and solving linear systems PhysRevLett.129.230504 . Additionally, VQAs have been extensively studied for simulating quantum dynamics, employing adaptive variational principles to optimize the time evolution of quantum states PRXQuantum.2.030307 . This approach is particularly important for simulating open quantum systems, where the interaction between system dynamics and environmental factors plays a crucial role PRXQuantum.5.020332 . Furthermore, VQAs have demonstrated their capability to achieve quantum-enhanced precision in multiparameter quantum metrology, even in the presence of noise Le2023 .
Quantum compilation, on the other hand, has gained significant interest due to its capacity to optimize quantum circuits through a training process that transforms target unitaries into trainable unitaries heya2018variational ; Khatri2019quantumassisted . This approach has been applied to various tasks, including gates optimization heya2018variational , quantum-assisted compiling Khatri2019quantumassisted , continuous-variable quantum learning PRXQuantum.2.040327 , quantum state tomography hai2023universal , and quantum objects simulation HAI2024101726 . For instance, a quantum object such as a quantum state can be prepared and its evolution can be simulated in a quantum circuit by using QC HAI2024101726 .
@@ I Introduction
The performance of QCs relies on the number of qubits and the circuit depth. The choice of ansatzes (trainable quantum circuits) is also crucial and must be carefully selected. Some entangled topologies have shown promise in solving quantum state preparation problems but require many resources due to their numerous layers and parameters hai2023variational .
Traditional QCs focus on compiling a single target heya2018variational ; Khatri2019quantumassisted ; PRXQuantum.2.040327 ; hai2023universal ; HAI2024101726 as mentioned above. However, many quantum systems require optimizing multiple factors at the same time, because independent optimization of each target can lead to suboptimal results in complex scenarios. For instance, simulating time-dependent quantum systems PhysRevA.99.042314 ; PRXQuantum.2.030307 ; Berry2024quantumalgorithm ; Mizuta2023optimalhamiltonian often requires balancing accuracy across various time intervals, while thermal-dependent systems Sagastizabal2021 ; PhysRevLett.123.220502 ; Zhu_2020 ; PhysRevApplied.16.054035 demand simultaneous optimization of thermal fluctuations and quantum coherence to maintain system stability. Similarly, in multimode tomography PhysRevA.90.053818 ; He2024 , optimizing multiple modes concurrently enhances the overall precision of quantum state estimation. These examples demonstrate that multi-target optimization is crucial in achieving optimal performance in quantum applications.
@@
This work presents a multi-target quantum compilation algorithm to address these demains by providing a unified framework for optimizing multiple objectives simultaneously. Therein, we compile multiple unitaries into a single trainable unitary through an optimization process, improving the performance of QC and applicability in practical scenarios. We explain the theoretical foundations of our algorithm, its structure, and using techniques like the genetic algorithm (GA) to enhance the multi-target QC process. We show the effectiveness of our algorithm through benchmarking and case studies, including the preparation of thermal states and the simulation of time-dependent Hamiltonians in comparison with other methods including Trotterlization Ikeda2023minimum ; ikeda2023trotter24 and adaptive variational quantum dynamics simulations (AVQDS) PRXQuantum.2.030307 . We also provide a demonstration for the variational quantum eigensolver. The results highlight the importance of multi-target compilation in advancing quantum computing technologies and the potential for further innovation in this field.
The novelty of this work includes:
  • It presents a quantum compilation algorithm for the first time that is adaptable for simulating multiple unitaries.
  • It introduces an enhanced quantum architecture search that utilizes a genetic algorithm in combination with a standard variational quantum algorithm.
The paper is structured as follows. In Section II, we present the multi-target quantum compilation with detail of its definitions, algorithms, and numerical benchmarking. Section III explores various applications, including thermal state preparation, time-dependent quantum dynamic simulations, and variational quantum eigensolvers. Finally, we conclude the paper in Section IV. Additional material can be found in A and B.
II Multi-target quantum compilation
@@
Quantum compilation (QC) Khatri2019quantumassisted ; hai2023variational can be regarded as a training process that transforms information from a target unitary $U$ to a trainable unitary $V(\bm{\theta})$, where $\bm{\theta}$ are trainable parameters. Hereafter, we introduce a compilation process to transform a set of $n$ targets into one trainable unitary (one single quantum circuit with $n$ different $\bm{\theta}$s), therefore, it is called multi-target quantum compilation.
II.1 Definition
@@
where $\bm{\theta}_{j}\in\big{\{}\theta_{j}^{(1)},\theta_{j}^{(2)},...,\theta_{j}^{(m)}\big{\}}$ is a set of $m$ trainable parameters and $\phi$ is a global phase. For each $U_{j}$, we need to maximize the kernel $\mathcal{K}_{U_{j},V}(\bm{\theta}_{j})=\frac{1}{u^{2}}\left|{\rm Tr}[U_{j}V^{\dagger}(\bm{\theta}_{j})]\right|^{2}$ where $u$ is the dimension of $U_{j}$. For maximizing all $\mathcal{K}_{U_{j},V}(\bm{\theta}_{j}),\forall U_{j}\in\mathbf{\mathcal{U}}$, we define a cost function $\mathcal{L}$, which is the average infidelity
$\mathcal{L}(\bm{\Theta},V)=1-\frac{1}{n}\sum_{j=1}^{n}\mathcal{K}_{U_{j},V}(\bm{\theta}_{j}),$ (2)
where $\bm{\Theta}=(\bm{\theta}_{1},\cdots,\bm{\theta}_{n})$ are trainable parameters, and $0\leq\mathcal{L}(\bm{\Theta},V)\leq 1$. The optimization process is to find $\bm{\Theta}^{*},V^{*}=\operatorname*{argmin}_{\{\bm{\Theta},V\}}\mathcal{L}(\bm{\Theta},V).$ Here, we emphasize that not only the trainable parameters are optimized but also the structure of ansatz $V$ is also optimized.
II.2 Quantum compilation algorithm
@@
where $G_{k}(\bm{\theta}_{k})$ is a sequence of parameterized single- and multi-qubit quantum gates, $L$ is the number of layers, $N$ the number of qubits. The design of $V(\bm{\theta})$ is not unique, and recent efforts to optimize it, including architecture search Du2022 ; Zhang_2021 ; PhysRevResearch.2.023074 ; https://doi.org/10.1002/qute.202100134 , adaptive variational quantum algorithm Grimsley2019 ; PRXQuantum.2.030307 , and genetic-based approach (GA) 934383 ; 8862255 ; katoch2021review ; Tandeitnik_2024 .
In this work, we combine GA with VQA to optimize multi-target quantum compilation. This approach optimizes both the parameters and the structure of quantum circuits (ansatzes), ensuring efficient compilation, practical computational capability, and low-depth circuits. The scheme is shown in Figure 1(a), and the flow chart is given in Figure 1(b). We first generate quantum circuits with random parameters and evaluates their cost functions. If the cost functions do not meet a threshold, we use VQA to optimize the parameters and re-evaluate. After VQA finishes, the best circuits are passed to the GA, which updates their structure through selection, crossover, and mutation for the next generation. Once GA completes, the best circuit is fed back into the VQA, which runs until the threshold is met or the cost function stabilizes.
Figure 1: Integrated Genetic Algorithm with Variational Quantum Algorithm (GA-VQA) for multi-target quantum compilation optimization. (a) The GA-VQA method uses a parameterized quantum circuit and a classical computer to evaluate and update both the circuit structure and its parameters. (b) The process is as follows: (i) Generate a set of circuits $\mathcal{V}$ and parameters $\bm{\Theta}$. (ii) Evaluate the cost function $\mathcal{L}(\bm{\Theta},V_{l})$ for each circuit $l\in[1,n_{\mathcal{V}}]$. (iii) If the threshold is not met, use VQA followed by GA to find the best circuit, checking the threshold after each iteration. (iv) If the threshold is still not reached, pass the best circuit $V^{*}$ back to VQA and repeat the process until the threshold is met. Here, $i$ and $g$ represent iteration indexes running up to $n_{\rm iter}$ and $n_{\rm gene}$, respectively.
Concretely, denote the set of generated quantum circuits as $\mathbf{\mathcal{V}}=\big{\{}V_{l};1\leq l\leq n_{\mathbf{\mathcal{V}}}\big{\}}$, where $n_{\mathbf{\mathcal{V}}}$ is the number of circuits. Each circuit $V_{l}\in\mathbf{\mathcal{V}}$ is created from a pool of quantum gates. For each circuit $V_{l}$, we evaluate the average infidelity via $\mathcal{L}(\bm{\Theta},V_{l})\ \forall 1\leq l\leq n_{\mathbf{\mathcal{V}}}.$ If $\mathcal{L}(\bm{\Theta},V_{l})\leq\rm threshold$, we designate $V_{l}$ to the best circuit, $V^{*}\leftarrow V_{l}$, and $\bm{\Theta}$ will be the optimal parameters, $\bm{\Theta}^{*}\leftarrow\bm{\Theta}$. If not, we run the VQA scheme with gradient descent to update $\bm{\Theta}$ for a certain number of iterations $n_{\rm iter}$. Then, we use GA with selection, crossover, and mutation to create the next generation of $\mathcal{V}$. See A for a detailed evolution of the GA. The GA scheme repeats until the threshold is met or the number of generations $n_{\rm gene}$ is reached. Finally, if $n_{\rm gene}$ is reached without meeting the threshold, we pass the best circuit $V^{*}$ to the VQA scheme again, and optimize $\bm{\Theta}$ until the threshold is met.
@@
Our scheme differs from the one-stage Du2022 and two-stage architecture search PhysRevResearch.2.023074 ; Ostaszewski2021structure . In those methods, all circuits are generated, their cost functions are evaluated as the low-depth and high number of gates as much as possible (gate density 10.1145/3550488 close to unity), and they are sorted from best to worst to select the top one. In our approach, each circuit $V_{l}\in\mathcal{V}$ is evaluated and compared with the best circuit (initially set as $V_{1}$). We thus do not require space to store all circuits. Although genetic algorithms have been suggested for quantum state tomography Creevey2023 , we believe they are also promising for different types of VQAs, including our proposed method for quantum compilation.
So far, we have observed that Ashhab et al. found that if a quantum circuit can perfectly implement one arbitrary target using random search, it can also implement any other targets with the same gate configuration. It only requires recalculating the single-qubit rotation parameters for each new target PhysRevA.106.022426 ; PhysRevA.109.052605 . This finding is similar to our discovery here using quantum compilation.
II.3 Numerical illustration
@@
For numerical evaluation, we follow Ref. Caro2023 and divide a random set $\mathcal{U}$ into a training set $\mathcal{U}_{\text{train}}=\{U_{1},U_{2},...,U_{\rm train}\}$ and a testing set $\mathcal{U}_{\text{test}}=\{U_{{\rm train+1}},U_{\rm train+2},...,U_{n}\}$. We then train the model on $\mathcal{U}_{\text{train}}$ using the cost function $\mathcal{L}(\bm{\Theta}_{\rm train},V_{l})$ as in Eq. 2. After training, we get the best circuit $V^{*}$ and evaluate the expected risk $\mathcal{R}$ over $\mathcal{U}_{\text{test}}$
@@
$\mathcal{R}(\bm{\Theta}_{\rm test})$ $=\frac{1}{4}\mathbb{E}_{U_{j}\;\in\;\mathcal{U}_{\text{test}}}\Big{\|}U_{j}|\bm{0}\rangle\langle\bm{0}|U_{j}^{\dagger}$
$\hskip 56.9055pt-[V^{*}]^{\dagger}(\bm{\theta}_{j})|\bm{0}\rangle\langle\bm{0}|V^{*}(\bm{\theta}_{j})\Big{\|}_{1}^{2},$ (4)
where $|\bm{0}\rangle\equiv|0\rangle^{\otimes N}$, with $N$ is the number of qubits.
@@
Figure 2: Numerical benchmarking for a set of Haar random $\mathcal{U}$. (a-d) Correlation between risk and fidelity for different number of generations $n_{\rm gene}$ and circuits per generation $n_{\mathcal{V}}$. (e) Plot of fidelity versus depth $d$ for different number of qubits $N$ ranging from 2 to 5. (f) The best quantum circuit generated by GA-VQA for $N=3$ with $d=9$. (g) The quantum circuit generated by Qiskit with $d=11$. (h) $|g_{2}g_{N}\rangle$ (2 layers) ansatz before transforming into single and two-qubit gates.
Figure  2(a-d) shows the correlation between risk $\mathcal{R}$ and fidelity $\mathcal{F}=1-\mathcal{L}$ for several $n_{\rm gene}$ and $n_{\mathcal{V}}$ with $N=3$. The GA-VQA method identifies a low-risk, high-fidelity ansatz for multi-target Haar random unitaries. Additionally, the factors $n_{\rm gene}$ and $n_{\mathcal{V}}$ have minimal impact on the results, allowing GA-VQA to run efficiently with minimal values, thus saving computational resources. The variation in some cases, such as the yellow lines in Figure  2(a-c), is a result of fluctuations in the optimization process for specific parameter settings. These fluctuations can arise from the randomized nature of the initial population in the genetic algorithm or the stochastic behavior of the variational quantum algorithm (VQA). Furthermore, the performance of different circuit architectures can vary significantly, leading to occasional deviations in the optimization trajectory.
@@
Figure  2(e) shows fidelity versus depth $d$ for 2-5 qubits. As expected, higher $N$ requires larger $d$ to achieve high fidelity. In the figure, circles mark the depth where $\mathcal{F}\geq 0.99$, with upper bounds generated by the standard Qiskit method. Figure  2(f) shows the best circuit generated by our GA-VQA compared to the Qiskit-generated circuit in Figure  2(g) and the $|g_{2}g_{N}\rangle$ ansatzhai2023variational in Figure  2(h) for $N=3$. The GA-VQA circuit has the shortest depth (9), compared to 11 and 38 for the others. Table 1 shows detailed comparisons for 2 to 5 qubits, demonstrating that the GA-VQA circuit provides high fidelity with the smallest depth.
@@
II.4 The complexity of GA-VQA
We discuss here the complexity of GA-VQA. Time complexity describes how the runtime of an algorithm changes as the input size increases, usually expressed using Big-O notation. It gives an estimate of the maximum runtime as the input grows. Common complexities include $\mathcal{O}(1)$ for constant time, $\mathcal{O}(z)$ for linear time, $\mathcal{O}(z^{2})$for quadratic time, and $\mathcal{O}(\log z)$ for logarithmic time. Here, $z$ is the input size. The time complexity of GA-VQA is defined as $\mathcal{O}(n\times n_{\rm gene}\times n_{\mathcal{V}}\times\mathcal{F})$ which depends on the number of targets $n$, the number of generations $n_{\rm gene}$, the number of circuits in each generation $n_{\mathcal{V}}$, and the fidelity $\mathcal{F}$. When the fidelity of circuits in one generation can be evaluated concurrent (parallel mode), the actual time complexity is $\mathcal{O}(n_{\rm gene}\times n_{\mathcal{V}}\times\mathcal{F})$, assuming that number of processing cores is larger than $n$.
The space complexity of an algorithm refers to the amount of memory or storage it requires relative to the size of the input. The space complexity in our case is $O(n\times n_{\mathcal{V}}\times\mathcal{F})$.
Since $n_{\rm gene},n_{\mathcal{V}}$ are hyper-parameters, the time and space complexities heavily depend on the cost function. For instance, for GA-VQA applied to multi-target compilation, the time complexity is approximately $\mathcal{O}(n_{\rm gene}\times n_{\mathcal{V}}\times n_{\text{iter}}\times 2^{N}\times d)$, where $d$ is the circuit depth and $n_{\text{iter}}$ is the number of optimization iterations. The space complexity for a state vector simulator is at least $\mathcal{O}(n\times n_{\mathcal{V}}\times 2^{N})$. For example, with 8 GB of RAM, GA-VQA can support up to 10 qubits with $n=5$ and $n_{\mathcal{V}}=4$, which results in approximately 6 GB used, calculated as follows: $5\times 4$ (for concurrent processes), $64\times 2\times(2^{10}+2^{20})$ (for a double-precision, 10-qubit quantum state-operator), multiplied by 2 for each buffer (state and circuit).
Currently, GA-VQA can manage unitaries that act on around 8-10 qubits with moderate circuit depths of approximately 40. More details can be found in B. However, this estimate may change with advancements in hardware capabilities, such as improved qubit coherence times and error rates. These factors are also the bottleneck for scaling the system size.
III Applications
III.1 Thermal state preparation (TSP)